# MathFusion: Enhancing Mathematic Problem-solving of LLM through Instruction Fusion

### v1

@@ Abstract
Large Language Models (LLMs) have shown impressive progress in mathematical reasoning. While data augmentation is promising to enhance mathematical problem-solving ability, current approaches are predominantly limited to instance-level modifications—such as rephrasing or generating syntactic variations—which fail to capture and leverage the intrinsic relational structures inherent in mathematical knowledge. Inspired by human learning processes, where mathematical proficiency develops through systematic exposure to interconnected concepts, we introduce MathFusion, a novel framework that enhances mathematical reasoning through cross-problem instruction synthesis. MathFusion implements this through three fusion strategies: (1) sequential fusion, which chains related problems to model solution dependencies; (2) parallel fusion, which combines analogous problems to reinforce conceptual understanding; and (3) conditional fusion, which creates context-aware selective problems to enhance reasoning flexibility. By applying these strategies, we generate a new dataset, MathFusionQA, followed by fine-tuning models (DeepSeekMath-7B, Mistral-7B, Llama3-8B) on it. Experimental results demonstrate that MathFusion achieves substantial improvements in mathematical reasoning while maintaining high data efficiency, boosting performance by 18.0 points in accuracy across diverse benchmarks while requiring only 45K additional synthetic instructions, representing a substantial improvement over traditional single-instruction approaches. Our datasets, models, and code are publicly available at https://github.com/QizhiPei/mathfusion.
MathFusion: Enhancing Mathematic Problem-solving of LLM
@@
Qizhi Pei1, Lijun Wu2∗, Zhuoshi Pan3, Yu Li2, Honglin Lin2, Chenlin Ming4, Xin Gao2, Conghui He2∗, Rui Yan1,5††thanks: Corresponding authors: Lijun Wu (wulijun@pjlab.org.cn), Conghui He (heconghui@pjlab.org.cn), and Rui Yan (ruiyan@ruc.edu.cn). 1Gaoling School of Artificial Intelligence, Renmin University of China 2Shanghai AI Laboratory 3Tsinghua University 4Shanghai Jiao Tong University 5School of Computer Science, Wuhan University {qizhipei,ruiyan}@ruc.edu.cn {wulijun,heconghui}@pjlab.org.cn
@@ 1 Introduction
Specialized mathematical LLMs have emerged to address the unique challenges of solving complex mathematical problems (Yang et al., 2024; Shao et al., 2024; Ying et al., 2024; team, 2024). Current approaches to enhance mathematical reasoning primarily focus on four paradigms: continued pre-training with math corpora (Yang et al., 2024; Shao et al., 2024), reinforcement learning (RL) from human or automated feedback (Luo et al., 2023; Lu et al., 2024), test-time compute scaling (Wang et al., 2024a; Kang et al., 2024; Guan et al., 2025; Xi et al., 2024), and supervised fine-tuning (SFT) using problem-solution pairs (Tang et al., 2024; Tong et al., 2024). Among these, SFT is the most widely adopted paradigm (Setlur et al., 2024) due to its simplicity. However, its effectiveness is often limited by the complexity and diversity of the mathematical training data (Luo et al., 2023) during SFT. To this end, data augmentation and synthesis have emerged as promising directions to enhance mathematical reasoning. For example, approaches such as MetaMath (Yu et al., 2024) and WizardMath (Luo et al., 2023) emphasize enhancing individual problems through rephrasing and difficulty variation.
@@
Motivated by the way human learners develop proficiency through systematic exposure to interconnected ideas (Komarudin et al., 2021), we propose MathFusion, a novel framework that enhances mathematical reasoning by fusing different mathematical problems. The key insight behind MathFusion is that the strategic combination of complementary mathematical instructions can unlock deeper reasoning capabilities. Specifically, by combining two existing problems, MathFusion synthesizes a new math problem that encapsulates the relational and compositional aspects of the original two problems. To achieve this, we introduce three distinct fusion strategies: (1) sequential fusion, which links related problems by chaining them together through shared variables to model solution dependencies; (2) parallel fusion, which integrates analogy problems to enhance conceptual comprehension and generate a novel problem that encapsulates their shared mathematical essence; and (3) conditional fusion, which generates selective problems based on specific context to promote flexible reasoning.
@@
Experimental results demonstrate that MathFusion enables LLMs to effectively capture the underlying relational structures of mathematical tasks, thereby enhancing their capacity to resolve complex, multi-step problems. Moreover, MathFusion yields considerable improvements in mathematical reasoning accuracy across both in-domain and challenging out-of-domain benchmarks, outperforming traditional single-instruction fine-tuning by 18.0 points in accuracy on average while incorporating only 45K additional synthetic instructions. Further integration with the state-of-the-art (SOTA) data augmentation method DART-Math (Tong et al., 2024) leads to additional improvements, surpassing it by 1.4 points in accuracy on average while utilizing less than one-third of the data employed by DART-Math. This highlights the complementary and orthogonal nature of our approach with existing data augmentation techniques.
@@ 2.1 Individual Data Augmentation for Math
Existing mathematical data augmentation methods primarily focus on two aspects: enhancing existing data and generating new data. Enhancing existing data typically involves modifying the problem or solution. For the problem, strategies include altering the level of complexity/difficulty (Luo et al., 2023), rephrasing the wording (Yu et al., 2024; Li et al., 2024b), and employing backward reasoning (Yu et al., 2024). For the solution, methods such as generating diverse and high-quality mathematical reasoning paths through multiple calls (Yu et al., 2024; Li et al., 2024b; Zhang et al., 2024; Tong et al., 2024), and incorporating reflection (Zhang et al., 2024) are commonly used. Generating new data typically involves creating new mathematical problems based on key mathematical concepts (Tang et al., 2024), seed datasets (Ding et al., 2024), specific example (Li et al., 2024a), and then using strong mathematical models (OpenAI et al., 2023; Shao et al., 2024) to generate corresponding solutions.
These methods, however, focus primarily on individual mathematical problems, overlooking the underlying relationships between different mathematical problems.
@@ 3 MathFusion
The overview of MathFusion is shown in Figure 2. Given two mathematical problems and from the original mathematical training set, MathFusion synthesizes a new mathematical problem by fusing these two problems. A simple example for and is shown in Example 3, and we show the corresponding for three fusion strategies in the following sections. More cases are shown in the Appendix F.
@@ 3.2 Sequential Fusion
The answer from solving serves as a part of the input to , thereby creating a chained dependency. A specific example of sequential fusion is shown in Example 3.2. The answer of the (the number of people transported by the boat) is used as the input for (the number of people in the first bus).
@@ 4.1 Experimental Setup
Training: We conduct standard instruction-tuning on our MathFusionQA. Following DART-Math (Tong et al., 2024), we conduct experiments on two categories of base models: 7B math-specialized base LLM, specifically DeepSeekMath-7B (Shao et al., 2024), and 7-8B general base LLMs, specifically Mistral-7B (Jiang et al., 2023) and Llama3-8B (Dubey et al., 2024). We fine-tune each base model with three fusion strategies—sequential, parallel, and conditional—each of which is the union of GSM8K, MATH, and the augmented set generated by the corresponding fusion strategy. Table 4 shows the statistics of the MathFusionQA collection. All models are trained for 3 epochs for simplicity. More details about the training setup are provided in Appendix B.2.
@@
Baselines: We mainly compare our MathFusion models with mathematical instruction-based models, which can be categorized into three groups: (1) Previous top-performing models, including MetaMath (Yu et al., 2024), WizardMath (Luo et al., 2023), RFT (rejection sampling fine-tuning) (Yuan et al., 2023; Tong et al., 2024), MMIQC (Liu et al., 2024), MathScale (Tang et al., 2024), DeepSeekMath-7B-Instruct (Shao et al., 2024), RefAug (Zhang et al., 2024), and DART-Math (Tong et al., 2024) (we report the Prop2Diff version); (2) Models instruction-tuned on the combination of GSM8K and MATH datasets (noted as ‘standard’ setting); (3) Models instruction-tuned on the sampled 60K version of previous top-performing methods to further evaluate the data efficiency of different mathematical data augmentation methods. Details about the sampling method are introduced in Appendix B.2.
@@ 4.3 Ablation Study
We further conduct an ablation study to investigate the contribution of each fusion strategy to the overall performance of combined fusion. The results over Llama3-8B on MATH and GSM8K are shown in Table 3, from which we observe that each fusion strategy contributes to the overall performance, with conditional fusion showing the least contribution, which aligns with Section 4.2.
@@ 5.3 Combination with Other Datasets
We further investigate the performance of MathFusion when combined with other data augmentation methods. Specifically, we downsample 30K-180K data from DART-Math-Hard (Tong et al., 2024), which is the SOTA method for mathematical data augmentation with 590K data. We combine the downsampled DART-Math-Hard with our MathFusionQA dataset and fine-tune Llama3-8B models on the combined dataset. The results are presented in Table 4. As the size of sampled data increases, the average performance of the models also increases, and reaches the peak when the size of the sampled data is 120K. Notably, by only using 90K data sampled from DART-Math-Hard (i.e., 150K samples in total), the resulting model achieves better performance than both DART-Math and MathFusion, yields SOTA average performance. These results show the potential of combining MathFusion with other data augmentation methods to further enhance the model’s performance. We think that the enhancement arises from the complementary and orthogonal nature of the two methods: our MathFusion emphasizes fusing mathematical problems to generate more challenging and diverse problems, while DART-Math focuses on existing difficult problems and primarily generates additional solutions for them.
@@ 6 Conclusion
We utilize strong GPT-4o-mini to generate fused problems and solutions, but the generated problems or solutions may still contain errors or ambiguities, which are hard to detect and verify. The quality of the generated problems and solutions is limited by the capabilities of the teacher LLM. We mainly explore the effectiveness of the three fusion strategies on problem pairs that are constructed by embedding similarity. The fusion of three or more problems and more effective ways to find similar problems, remain unexplored.
@@ Appendix A Prompts
We show the prompts used for Sequential Fusion in Prompt E, Parallel Fusion in Prompt E, and Conditional Fusion in Prompt E. We also provide the problem evaluation prompts in Prompt E, which is partially derived from WizardMath (Luo et al., 2023). We use LangGPT (Wang et al., 2024b) to format prompts in Markdown and polish them.
@@ C.1 Fused Probelm Types
| Standard | 15K | 17.5 | 65.4 | 12.9 | 21.6 | 4.7 | 10.9 | 22.2 |
| Standard + GPT Rewritten | 30K | 22.8 | 75.4 | 11.8 | 15.7 | 5.5 | 9.6 | 23.5 |
| MathFusion (Sequential) | 30K | 38.8 | 77.9 | 25.1 | 42.0 | 12.6 | 17.0 | 35.6 |
| MathFusion (Parallel) | 30K | 38.1 | 75.4 | 25.5 | 41.9 | 11.9 | 18.9 | 35.3 |
| MathFusion (Conditional) | 30K | 34.7 | 76.9 | 21.2 | 27.4 | 11.9 | 15.5 | 31.3 |
| MathFusion | 60K | 46.5 | 79.2 | 27.9 | 43.4 | 17.2 | 20.0 | 39.0 |
@@ C.2 Fused Error Analysis
In practice, we find that some fused problems are unreasonable or ambiguous, which are shown in Section F. The reason may be that some problems are not suitable for fusion or the limited capacity of the model for generating fused problems. To verify the correctness of the fused problems and their influence on the model’s performance, we conduct an error analysis on the fused problems. Specifically, borrowing the idea from rejection sampling (Yuan et al., 2023), we use GPT-4o-mini to verify the correctness and completeness of the fused problems. The corresponding evaluation prompt is shown in Section A. For each identified unreasonable problem, we adjust the temperature to 1.0 to enhance the diversity of generation, and re-generate the problems five times using the corresponding fusion strategy. If none of the five generated problems is reasonable, we consider the fusion to be unreasonable and discard it. Finally, 5.6% of the fused problems are identified as unreasonable, and the remaining reasonable problems are added to the dataset. The average performance of Llama3-8B fine-tuned only on the filtered MathFusionQA is 39.1, which is similar to the performance of the model fine-tuned on the original MathFusionQA (39.0), indicating that the unreasonable problems have little impact on the model’s performance.

### v2

# MathFusion: Enhancing Mathematical Problem-solving of LLM through Instruction Fusion

@@ Abstract
Large Language Models (LLMs) have shown impressive progress in mathematical reasoning. While data augmentation is promising to enhance mathematical problem-solving ability, current approaches are predominantly limited to instance-level modifications—such as rephrasing or generating syntactic variations—which fail to capture and leverage the intrinsic relational structures inherent in mathematical knowledge. Inspired by human learning processes, where mathematical proficiency develops through systematic exposure to interconnected concepts, we introduce MathFusion, a novel framework that enhances mathematical reasoning through cross-problem instruction synthesis. MathFusion implements this through three fusion strategies: (1) sequential fusion, which chains related problems to model solution dependencies; (2) parallel fusion, which combines analogous problems to reinforce conceptual understanding; and (3) conditional fusion, which creates context-aware selective problems to enhance reasoning flexibility. By applying these strategies, we generate a new dataset, MathFusionQA, followed by fine-tuning models (DeepSeekMath-7B, Mistral-7B, Llama3-8B) on it. Experimental results demonstrate that MathFusion achieves substantial improvements in mathematical reasoning while maintaining high data efficiency, boosting performance by 18.0 points in accuracy across diverse benchmarks while requiring only 45K additional synthetic instructions, representing a substantial improvement over traditional single-instruction approaches.
MathFusion: Enhancing Mathematical Problem-solving of LLM
@@
Qizhi Pei1,2,
Lijun Wu2∗,
Zhuoshi Pan2,3,
Yu Li2,
Honglin Lin2,
Chenlin Ming2,4,
Xin Gao2,
Conghui He2∗,
Rui Yan1,5,6††thanks: Corresponding authors: Lijun Wu (wulijun@pjlab.org.cn), Conghui He (heconghui@pjlab.org.cn), and Rui Yan (ruiyan@ruc.edu.cn).
1Gaoling School of Artificial Intelligence, Renmin University of China
2Shanghai Artificial Intelligence Laboratory 3Tsinghua University 4Shanghai Jiao Tong University
5Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE
6School of Artificial Intelligence, Wuhan University
{qizhipei,ruiyan}@ruc.edu.cn {wulijun,heconghui}@pjlab.org.cn
https://huggingface.co/collections/QizhiPei/MathFusion
https://github.com/QizhiPei/MathFusion
@@ 1 Introduction
Specialized mathematical LLMs have emerged to address the unique challenges of solving complex mathematical problems (Yang et al., 2024; Shao et al., 2024; Ying et al., 2024; team, 2024). Current approaches to enhance mathematical reasoning primarily focus on four paradigms: continued pre-training with math corpora (Yang et al., 2024; Shao et al., 2024), reinforcement learning (RL) from human or automated feedback (Luo et al., 2023; Lu et al., 2024), test-time compute scaling (Wang et al., 2024a; Kang et al., 2024; Guan et al., 2025; Xi et al., 2024), and supervised fine-tuning (SFT) using problem-solution pairs (Tang et al., 2024; Tong et al., 2024; Lin et al., 2025). Among these, SFT is the most widely adopted paradigm (Setlur et al., 2024) due to its simplicity. However, its effectiveness is often limited by the complexity and diversity of the mathematical training data (Luo et al., 2023) during SFT. To this end, data augmentation and synthesis have emerged as promising directions to enhance mathematical reasoning. For example, approaches such as MetaMath (Yu et al., 2024) and WizardMath (Luo et al., 2023) emphasize enhancing individual problems through rephrasing and difficulty variation.
@@
Motivated by the way human learners develop proficiency through systematic exposure to interconnected ideas (Komarudin et al., 2021), we propose MathFusion, a novel framework that enhances mathematical reasoning by fusing different mathematical problems. The key insight behind MathFusion is that the strategic combination of complementary mathematical instructions can unlock deeper reasoning capabilities. Specifically, by combining two existing problems, MathFusion synthesizes a new math problem that encapsulates the relational and compositional aspects of the original two problems. To achieve this, we introduce three distinct fusion strategies: (1) sequential fusion, which links related problems by chaining them together through shared variables to model solution dependencies; (2) parallel fusion, which integrates analogy problems to enhance conceptual comprehension and generates a novel problem that encapsulates their shared mathematical essence; and (3) conditional fusion, which generates selective problems based on specific context to promote flexible reasoning.
@@
Experimental results demonstrate that MathFusion enables LLMs to effectively capture the underlying relational structures of mathematical tasks, thereby enhancing their capacity to resolve complex, multi-step problems. Moreover, MathFusion yields considerable improvements in mathematical reasoning accuracy across both in-domain and challenging out-of-domain benchmarks, outperforming traditional single-instruction fine-tuning by 18.0 points in accuracy on average while incorporating only 45K additional synthetic instructions. Further integration with the state-of-the-art (SOTA) data augmentation method DART-Math (Tong et al., 2024) and scaling MathFusion to larger size lead to additional improvements, surpassing DART-Math in accuracy on average while utilizing less than one-third of its data. This highlights the complementary nature and scalability of our approach.
@@ 2.1 Individual Data Augmentation for Math
Existing mathematical data augmentation methods primarily focus on two aspects: enhancing existing data and generating new data. Enhancing existing data typically involves modifying the problem or solution. For the problem, strategies include altering the level of complexity/difficulty (Luo et al., 2023), rephrasing the wording (Yu et al., 2024; Li et al., 2024b), and employing backward reasoning (Yu et al., 2024). For the solution, methods such as generating diverse and high-quality mathematical reasoning paths through multiple calls (Yu et al., 2024; Li et al., 2024b; Zhang et al., 2024; Tong et al., 2024), and incorporating reflection (Zhang et al., 2024; Pan et al., 2025) are commonly used. Generating new data typically involves creating new mathematical problems based on key mathematical concepts (Tang et al., 2024), seed datasets (Ding et al., 2024), specific example (Li et al., 2024a), and then using strong mathematical models (OpenAI et al., 2023; Shao et al., 2024) to generate corresponding solutions. These methods, however, focus primarily on individual mathematical problems, overlooking the underlying relationships between different problems.
@@ 3 MathFusion
The overview of MathFusion is shown in Figure 2. Given two mathematical problems and from the original mathematical training set, MathFusion synthesizes a new mathematical problem by fusing these two problems. A simple example for and is shown in Example 3, and we show the corresponding for three fusion strategies in the following sections. More cases are shown in the Appendix G.
@@ 3.2 Sequential Fusion
The answer from solving serves as a part of the input to , thereby creating a chained dependency. A specific example of sequential fusion is shown in Example 3.2. The answer of (the number of people transported by the boat) is used as the input for (the number of people in the first bus).
@@ 4.1 Experimental Setup
Training: We conduct standard instruction-tuning on our MathFusionQA. Following DART-Math (Tong et al., 2024), we conduct experiments on two categories of base models: 7B math-specialized base LLM, specifically DeepSeekMath-7B (Shao et al., 2024), and 7-8B general base LLMs, specifically Mistral-7B (Jiang et al., 2023) and Llama3-8B (Dubey et al., 2024). Each base model is fine-tuned using three distinct fusion strategies: sequential, parallel, and conditional. For each strategy, the fine-tuning dataset (30K samples) comprises the union of the GSM8K, MATH datasets, and an augmented set generated by the respective fusion strategy. The MathFusionQA dataset (60K samples) is formed by the union of all these sub-datasets. Table 4 shows the statistics of the MathFusionQA collection. To demonstrate the scaling ability of MathFusion, we also enlarge MathFusionQA dataset by including top-2 to top-4 nearest neighbors, resulting in a total of 15K + 4 × (3 × 15K) = 195K examples. All models are trained for 3 epochs. More details about the training setup are provided in Appendix B.2.
@@
Baselines: We mainly compare our MathFusion models with mathematical instruction-based models, which can be categorized into three groups: (1) Previous top-performing models, including MetaMath (Yu et al., 2024), WizardMath (Luo et al., 2023), RFT (rejection sampling fine-tuning from DART-Math) (Yuan et al., 2023; Tong et al., 2024), MMIQC (Liu et al., 2024), MathScale (Tang et al., 2024), DeepSeekMath-7B-Instruct (Shao et al., 2024), RefAug (Zhang et al., 2024), and DART-Math (Tong et al., 2024) (we report the Prop2Diff version as it generally performs better than the Uniform version); (2) Models instruction-tuned on the combination of GSM8K and MATH datasets (noted as “Standard” setting); (3) Models instruction-tuned on the sampled 60K version of previous top-performing methods to further evaluate the data efficiency of different mathematical data augmentation methods. Details about the sampling method are introduced in Appendix B.2.
@@ 4.2 Main Results
Finding 5: MathFusion exhibits strong scalability and outperforms larger-scale baselines with fewer samples. The results on DeepSeekMath-7B (the strongest base model for math) in Table 2 reveal that scaling MathFusion from 60K to 195K samples leads to consistent performance gains across all evaluation benchmarks. Notably, MathFusion-DSMath-7B (195K) surpasses DeepSeekMath-7B-DART-Math (590K) in average accuracy (49.9 vs. 49.4), despite using only one-third of the training data. This illustrates the high scalability and data efficiency of MathFusion. The substantial performance gains on challenging benchmarks such as MATH (+4.6), DeepMind-Mathematics (+7.5), and OlympiadBench-Math (+3.8) underscore the method’s capability to generalize well under increased data volume. These findings demonstrate that MathFusion benefits substantially from larger synthetic training sets and can outperform significantly larger instruction-tuned models with less data.
@@ 4.3 Ablation Study
We further conduct an ablation study to investigate the contribution of each fusion strategy to the overall performance of combined fusion. The results over Llama3-8B on MATH and GSM8K are shown in Table 3, from which we observe that each fusion strategy contributes to the overall performance, with conditional fusion showing the least contribution, which aligns with Section 4.2. We further ablate on choice of teacher model for solution generation in Appendix D.
@@ 5 Analysis
We analyze the difficulty of the fused problem in Section 5.1, the relationship between augmented data size and performance in Section 5.2, the combination of MathFusionQA with other datasets in Section 5.3, and the diversity of fused problem in Section 5.4.
@@ 5.3 Combination with Other Datasets
We further investigate the performance of MathFusion when combined with other data augmentation methods. Specifically, we downsample 30K-180K data from DART-Math-Hard (Tong et al., 2024), which is the SOTA method for mathematical data augmentation with 590K data. We combine the downsampled DART-Math-Hard with our MathFusionQA dataset and fine-tune Llama3-8B models on the combined dataset. The results are presented in Figure 4. As the size of sampled data increases, the average performance of the models also increases, and reaches the peak when the size of the sampled data is 120K. Notably, by only using 90K data sampled from DART-Math-Hard (i.e., 150K samples in total), the resulting model achieves better performance than both DART-Math and MathFusion, yields SOTA average performance. These results show the potential of combining MathFusion with other data augmentation methods to further enhance the model’s performance. We think that the enhancement arises from the complementary and orthogonal nature of the two methods: our MathFusion emphasizes fusing mathematical problems to generate more challenging and diverse problems, while DART-Math focuses on existing difficult problems and primarily generates additional solutions for them.
@@ 6 Conclusion
We utilize GPT-4o-mini to generate fused problems and solutions, but the generated problems or solutions may still contain errors or ambiguities, which are hard to detect and verify. The quality of the generated problems and solutions is limited by the capabilities of the teacher LLM. Stronger teacher model, like DeepSeek-R1 and Qwen3, are underexplored. We mainly explore the effectiveness of the three fusion strategies on problem pairs that are constructed by embedding similarity. The fusion of three or more problems and more effective ways to find similar problems, remain underexplored. The released MathFusionQA dataset currently contains only 60K examples, and scaling to millions of examples remains underexplored.
Acknowledgements
This work is supported by Beijing Outstanding Young Scientist Program NO. BJJWZYJH012019100020098 and by National Key R&D Program of China (2022ZD0160201). This work is also supported by the Public Computing Cloud, Renmin University of China and by fund for building world-class universities (disciplines) of Renmin University of China. Qizhi Pei is supported by the Outstanding Innovative Talents Cultivation Funded Programs 2023 of Renmin University of China. Qizhi Pei is an intern at Shanghai Artificial Intelligence Laboratory.
@@ Appendix A Prompts
We show the prompts used for Sequential Fusion in Prompt F, Parallel Fusion in Prompt F, and Conditional Fusion in Prompt F. We also provide the problem evaluation prompts in Prompt F, which is partially derived from WizardMath (Luo et al., 2023). We use LangGPT (Wang et al., 2024b) to format prompts in Markdown and polish them.
@@ C.1 Fused Probelm Types
| Standard | 15K | 12.4 | 60.3 | 8.4 | 17.0 | 2.2 | 7.6 | 18.0 |
| GPT Rewritten | 15K | 20.1 | 70.3 | 9.1 | 13.9 | 2.8 | 8.9 | 20.9 |
| Mosaic-IT | 15K | 11.7 | 40.9 | 7.4 | 9.2 | 2.7 | 9.9 | 13.6 |
| Mosaic-IT + Original GSM8K and MATH | 30K | 11.0 | 54.7 | 6.9 | 9.8 | 1.9 | 9.5 | 15.6 |
| DART-Math | 60K | 34.1 | 77.2 | 23.4 | 36.0 | 8.7 | 18.2 | 32.9 |
| MathFusion (Sequential) | 30K | 32.7 | 73.9 | 18.9 | 29.3 | 9.3 | 15.5 | 29.9 |
| MathFusion (Parallel) | 30K | 30.9 | 75.1 | 20.9 | 26.5 | 11.0 | 15.2 | 29.9 |
| MathFusion (Conditional) | 30K | 26.3 | 73.0 | 15.6 | 21.4 | 7.3 | 12.8 | 26.1 |
| MathFusion (DeepSeekMath-7B-RL) | 60K | 42.0 | 78.1 | 24.0 | 36.5 | 13.0 | 13.8 | 34.6 |
| MathFusion | 60K | 41.6 | 79.8 | 24.3 | 39.2 | 13.6 | 18.1 | 36.1 |
@@ C.2 Fused Error Analysis
In practice, we find that some fused problems are unreasonable or ambiguous, which are shown in Section G. The reason may be that some problems are not suitable for fusion or the limited capacity of the model for generating fused problems. To verify the correctness of the fused problems and their influence on the model’s performance, we conduct an error analysis on the fused problems. Specifically, borrowing the idea from rejection sampling (Yuan et al., 2023), we use GPT-4o-mini to verify the correctness and completeness of the fused problems. The corresponding evaluation prompt is shown in Section A. For each identified unreasonable problem, we adjust the temperature to 1.0 to enhance the diversity of generation, and re-generate the problems five times using the corresponding fusion strategy. If none of the five generated problems is reasonable, we consider the fusion to be unreasonable and discard it. Finally, 5.6% of the fused problems are identified as unreasonable, and the remaining reasonable problems are added to the dataset. The average performance of Llama3-8B fine-tuned only on the filtered MathFusionQA is 39.1, which is similar to the performance of the model fine-tuned on the original MathFusionQA (39.0), indicating that the unreasonable problems have little impact on the model’s performance. This result aligns with the findings in OpenMathInstruct-2 (Toshniwal et al., ), indicating models exhibit some robustness to low-quality data in SFT.